# MetaCLBench: Meta Continual Learning Benchmark on Resource-Constrained Edge Devices

### v1

@@ 4.2 Analysis and Discussion
Effectiveness of Pre-training. We examine the effectiveness of pre-training on the performance of Meta-CL methods. Our evaluation across five datasets without pre-training reveals varying performance levels: while the two image datasets and GSCv2 demonstrated satisfactory accuracy, Urbansound and ESC50 showed suboptimal performance. Following the approach used in advanced models, we implemented pre-training before meta-training for each method. As shown in Figure 3a, the incorporation of pre-training significantly enhanced model performance. Notably, Meta-CL methods achieved an average performance improvement of 13.98% on the Urbansound and ESC50 datasets after pre-training. Our analysis revealed that CNNs exhibited considerable performance degradation when meta-training was conducted without pre-training. Given that advanced models inherently contain pre-trained weights, additional pre-training did not yield substantial accuracy improvements. These findings establish pre-training as a crucial step in optimizing model accuracy and demonstrate that the sequential implementation of pre-training followed by meta-training is essential for achieving superior performance across all methods.
@@
The Number of Replay Epochs. We investigated the relationship between replay frequency and accuracy, acknowledging that increased replays correspond to higher latency and energy consumption. As illustrated in Figure 3b, Oracle OML requires more than five replays to achieve high accuracy levels, resulting in increased energy consumption during training. In contrast, Latent OML demonstrates superior efficiency, achieving comparable accuracy with only one or two replays, thus reducing system overhead. Within the same dataset, Latent OML’s performance after five replays approximates that of Oracle OML after ten replays.
(a) With and Without pre-training on CNN
(b) Number of replay epochs
Figure 3: The analysis of Meta-CL Methods for the audio datasets
@@
Multi Compression Module. In CNN architectures, ANML with AIM demonstrates substantial accuracy improvements (from 26.0% to 34.6% on CIFAR100), albeit at the cost of increased memory requirements (from 39.69MB to 1,093MB). LifeLearner presents an optimal balance between accuracy and memory utilization in resource-constrained environments. The integration of AIM, while beneficial for accuracy, introduces significant computational overhead, particularly in Raw ANML implementations (11,424s latency, 52.5KJ energy consumption on CIFAR100). Oracle ANML provides a balanced compromise between performance enhancement and resource efficiency.
@@
Vision Transformer (ViT) and YAMNet architectures exhibit analogous performance-resource tradeoffs. ViT implementations with OML and AIM show improved accuracy but incur substantial computational costs (22,968s latency, 103KJ energy consumption on CIFAR100). Similarly, YAMNet with AIM achieves superior accuracy (from 42.9% to 71.0% on GSCv2) while requiring increased memory utilization (from 10.16MB to 608.2MB). Pretrained YAMNet models demonstrate superior resource efficiency metrics.
5 Practical Guidelines
@@
Our extensive empirical analysis provides evidence-based guidelines for researchers and practitioners implementing Meta-CL on edge devices. Although this investigation focuses on specific Meta-CL methods and may not encompass all learning approaches, our results and findings show consistent patterns across different architectures and datasets, which lead to the following key guidelines.
Resource-Aware Method Selection. The choice of Meta-CL implementation should be guided by available system resources. For systems with adequate storage capacity, LifeLearner or Latent OML implementations provide optimal accuracy across datasets. Resource-constrained environments benefit from Raw ANML or Raw OML methods combined with hardware-friendly optimizations. In severely limited environments, efficient integration of the quantization and compression module of LifeLearner with advanced models (YAMNet and ViT) would offer a viable compromise between computational efficiency and storage requirements.
@@
Edge Deployment Considerations. Latent methods consistently demonstrate superior performance across image and audio datasets, balancing accuracy with resource efficiency. This characteristic positions these methods as particularly advantageous for IoT applications, facilitating real-time category integration, efficient personalized processing, and privacy-preserving local computation without additional communication overhead.
6 Conclusion
In this paper, we investigate how to solve the CF problem based on five representative image and audio datasets (i.e., CIFAR100, MiniImageNet, GSCV2, UrbanSound8K, and ESC50) using three different network architectures with different complexities using six well-known Meta-CL methods. We also created a comprehensive Meta-CL benchmark framework on an edge device to assess the system overheads occurring on the device and examine the trade-offs between performance, computational costs, memory usage, and storage of different Meta-CL approaches. We first found that the combination of pre-training plus meta-training plus evaluation maximizes performance. Advanced methods (YAMNet and ViT) already include pre-trained, so the extra pre-training will not affect the results much.
Supplementary Material
MetaCLBench: Meta Continual Learning Benchmark on Resource-Constrained Edge Devices

### v2

@@ IV-B Analysis and Discussion
Effectiveness of Pre-training. We examine the effectiveness of pre-training on the performance of Meta-CL methods. Our evaluation across five datasets without pre-training reveals varying performance levels: while the two image datasets and GSCv2 demonstrated satisfactory accuracy, UrbanSound8K and ESC-50 showed suboptimal performance. Following the approach used in advanced models, we implemented pre-training before meta-training for each method. As shown in Figure 3, the incorporation of pre-training significantly enhanced model performance. Notably, Meta-CL methods achieved an average performance improvement of 13.98% on the UrbanSound8K and ESC-50 datasets after pre-training. Our analysis revealed that CNNs exhibited considerable performance degradation when meta-training was conducted without pre-training. Pre-training plays a significant role in enhancing accuracy, particularly for ANML and OML, where accuracy without pre-training is a mere 5%. Advanced architectures such as ViT and YAMNet intrinsically incorporate pre-training, obviating the need for ablation study of pre-training with them.
@@
The Number of Replay Epochs. We investigated the relationship between replay frequency and accuracy, acknowledging that increased replays correspond to higher latency and energy consumption. Recognizing that an increased number of replays correlates with heightened latency and energy usage, we explored the association between replay frequency and accuracy. Figure  4a indicates that Oracle OML necessitates more than five replays to attain elevated accuracy levels, thereby escalating energy expenditure throughout the training phase. Conversely, Latent OML demonstrates superior performance, requiring merely one or two replays to reach comparable accuracy, thereby diminishing system overhead. In the case of the same dataset, the efficiency of Latent OML after five replays approximates that of Oracle OML following ten replays. The same observation is evident in the CNN architecture depicted in Figure  4c; here, one or two iterations of the Latent method attain the same level of accuracy as the Oracle method after five iterations. Additionally, we note that exceeding five iterations does not significantly enhance accuracy. Figures 4b and 4d illustrate that the accuracy of Raw OML within the advanced methods surpasses that of the previously tested Latent approach.
@@
Analysis of AIM and Compression Module. In CNN architectures, ANML with AIM demonstrates substantial accuracy improvements (from 26.0% to 34.6% on CIFAR-100), albeit at the cost of increased memory requirements (from 39.69 MB to 1,093 MB). The integration of AIM, while beneficial for accuracy, introduces significant computational overhead. LifeLearner presents an optimal balance between accuracy and memory utilization in resource-constrained IoT devices via efficient compression module for latent replays.
@@
Vision Transformer (ViT) and YAMNet architectures exhibit analogous performance-resource tradeoffs. ViT implementations with OML and AIM show improved accuracy but incur substantial computational costs (22,968s latency, 103 kJ energy consumption on CIFAR-100). Similarly, YAMNet with AIM achieves superior accuracy (from 42.9% to 71.0% on GSCv2) while requiring increased memory utilization (from 10.16 MB to 608.2 MB). Pre-trained YAMNet models demonstrate superior resource efficiency metrics.
IV-C Differences between image and audio datasets
IoT deployments span both image sensors (cameras, visual inspection systems) and audio sensors (microphones for voice interfaces, acoustic monitoring). Understanding modality-specific Meta-CL behavior is essential for system design.
First, the temporal nature of audio data creates distinct feature dependencies compared to image data. Our experimental results demonstrate that ANML achieves 59.6% accuracy on UrbanSound8K versus only 32.7% on MiniImageNet, a disparity stemming from audio’s sequential characteristics being better suited for neural networks to extract temporal patterns.
Also, audio data requires specialized preprocessing steps (e.g., spectrogram conversion or MFCC feature extraction) that introduce extra complexity in Meta-CL systems. Architectures optimized for audio like YAMNet exhibit different memory-energy tradeoffs compared to visual architectures like ViT, with the former showing better energy efficiency on audio samples (5.4 kJ for UrbanSound8K vs. 11.2 kJ for MiniImageNet).
Finally, we find that audio Meta-CL is more sensitive during the pre-training phase, as in Figure 3, where pre-training improved average performance by 13.98% for UrbanSound8K and ESC-50 datasets, while the improvement for image datasets was substantially smaller. This indicates that knowledge transfer in the audio domain requires stronger domain-specific foundations to work effectively in Meta-CL settings.
(a) Oracle & Latent (YAMNet)
(b) Raw & Latent (YAMNet)
(c) Oracle & Latent (CNN)
(d) Raw & Oracle & Latent (ViT)
V Practical Guidelines
@@
Our extensive empirical analysis provides evidence-based guidelines for researchers and practitioners implementing Meta-CL on edge devices. While our investigation focuses on specific Meta-CL methods and may not encompass all learning approaches, our results and findings show consistent patterns across different architectures and datasets, which lead to the following key guidelines.
Resource-Aware Method Selection. Device memory during runtime (not nominal RAM) is the primary constraint determining method viability. For devices with less than 1 GB available RAM, AIM-enhanced methods should be excluded due to OOM failures observed in our experiments. LifeLearner and Latent OML provide the best balance of accuracy and resource efficiency across memory-constrained environments. For devices with sufficient memory ($\geq$2 GB available), AIM-enhanced methods become viable when accuracy is paramount.
Architecture Selection. Direct comparison across architectures is limited because CNN uses ANML-based methods while ViT and YAMNet use OML-based methods. Nevertheless, we observe that advanced architectures (ViT and YAMNet) do not necessarily guarantee better Meta-CL performance over CNN. For image tasks, results are dataset-dependent: CNN with Oracle ANML outperforms ViT with Oracle OML on CIFAR-100 (44.5% vs. 36.1%) but underperforms on MiniImageNet (43.8% vs. 51.2%). For audio tasks, YAMNet tends to achieve higher accuracy than CNN while requiring less memory. These findings challenge conventional assumptions about model complexity.
Pre-training Strategy. Pre-training before meta-training is essential, particularly for audio domains where we observed 14% average accuracy improvement (Figure 3). As detailed in Section III-F, for CNNs, we pre-train on source classes until the validation loss converges. For YAMNet and ViT, we leverage publicly available pre-trained weights directly.
@@
Edge Deployment Considerations. Latent methods such as LifeLearner for CNN and Latent OML for ViT/YAMNet consistently demonstrate strong performance across image and audio datasets, balancing accuracy with resource efficiency. These methods are particularly advantageous for battery-powered IoT applications, enabling real-time category integration, efficient personalised processing, and privacy-preserving local computation without additional communication overhead.